The cost per equivalent units of direct materials and conversion in the Bottling Department of Beverages on Jolt Company is $0.60 and $0.10, respectively. The equivalent units to be assigned costs are as follows: Direct Materials Conversion Inventory in process, beginning of period 0 2,540 Started and completed during the period 53,330 53,330 Transferred out of Bottling (completed) 53,330 55,870 Inventory in process, end of period 3,570 2,000 Total units to be assigned costs 56,900 57,870 The beginning work in process inventory had a cost of $3,030. Determine the cost of completed and transferred out production, and the ending work in process inventory. Completed and transferred out production $ Inventory in process, ending $
Answer #1
1
Beginning work in process inventory 3030
To complete Beginning work in process inventory 254 =2540*0.1
Units started and completed 37331 =53330*(0.6+0.1)
Completed and transferred out production 40615
2
Inventory in process, ending 2342 =(3570*0.6)+(2000*0.1)